OpenStudy (anonymous):

Illustrate the use of the quadratic formula to solve the same equation: x² + 6x + 9 = 25

OpenStudy (texaschic101):

x^2 + 6x + 9 = 25 we want this in this form : ax^2 + bx + c = 0 so we will put it in that form x^2 + 6x + 9 = 25 --- subtract 25 from both sides x^2 + 6x + 9 - 25 = 0 -- combine like terms x^2 + 6x - 16 = 0 now we will use quadratic formula: x = (-b (+/-) sq rt b^2 - 4ac) / (2a) x^2 + 6x - 16 = 0 a = 1, b = 6, c = -16 now we sub into the quadratic formula.. x = (-6 (+/-) sq rt 6^2 - 4(1)(-16) / (2(1)) x = (-6 (+/-) sq rt 36 + 64) / 2 x = (-6 (+/-) sq rt 100) / 2 x = (-6 (+/-) 10) / 2 x = (-6 + 10) / 2 x = 4/2 x = 2 and x = (-6 - 10)/2 x = -16/2 x = -8 so your solution is : x = 2 and x = -8

OpenStudy (texaschic101):

read this..it might help you.. http://www.purplemath.com/modules/quadform.htm